### v1

@@ 3 Query-Restricted Adversaries
and thus $\operatorname{PLD}(\Gamma)\succeq L\otimes\operatorname{PLD}(\Gamma_{L})$, which completes the proof since $L$ was arbitrary. ∎
Intuitively, this result says that a worst-case adversary plays a move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma_{L})$ to ensure that the overall process “traces out” the privacy curve of $\operatorname{Adv}(\Gamma_{L})$.
@@ 4 Characterizing the Natural Filter
Our goal is to study when the natural filter Algorithm˜1 with budget $B$ and queries in $\mathcal{L}$ comes at no additional privacy cost, i.e. the mechanism induced by the interaction between the filter and an adversarial analyst itself has privacy bounded by $B$. In this case, we say that the natural filter is free. We will write $L\ominus L^{\prime}:=\sup\{L^{\prime\prime}:L^{\prime}\oplus L^{\prime\prime}\preceq L\}$ so that $L^{\prime\prime}\preceq L\ominus L^{\prime}$ exactly when $L^{\prime}\oplus L^{\prime\prime}\preceq L$. We also write $\mathcal{L}_{\preceq L}:=\{L^{\prime}\in\mathcal{L}:L^{\prime}\preceq L\}$ to denote the queries in $\mathcal{L}$ dominated by $L$. For simplicity, we will assume that $\text{Id}\in\mathcal{L}$.
@@
0: Privacy budget $L$, adversary (or analyst) $\mathcal{A}$, family of allowed PLDs $\mathcal{L}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that has PLD $L_{i}\in\mathcal{L}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $L_{1}\oplus\dots\oplus L_{i}\preceq L$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
Now, notice that the natural filter Algorithm˜1 with budget $L$ and $k$ queries in $\mathcal{L}$ is free exactly when $\operatorname{PLD}(\operatorname{Filter}_{\mathcal{L},B,k})\preceq B$ where
$\operatorname{Filter}_{\mathcal{L},B,k}(L_{1},\dots,L_{k^{\prime}}):=\mathcal{L}_{\preceq B\ominus L_{1}\dots\ominus L_{k^{\prime}}}=\{L_{k^{\prime}+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{k^{\prime}}\oplus L_{k^{\prime}+1}\preceq B\}$
for $0\leq k^{\prime}<k$. In particular, by applying Lemma˜1 to the rule $\operatorname{Filter}_{\mathcal{L},B,k}$ we immediately get the following characterization.
@@
Let $\mathcal{L}$ be a family of PLDs, let $B$ be a PLD, and let $k$ be a positive integer. For any $L_{1},\dots,L_{k^{\prime}}\in\mathcal{L}$ ($0\leq k^{\prime}<k$) such that $L_{1}\oplus\dots\oplus L_{k^{\prime}}\preceq B$, let
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):=\sup\{L_{k^{\prime}+1}\oplus V_{k-(k^{\prime}+1)}(L_{1},\dots,L_{k^{\prime}+1}):L_{k^{\prime}+1}\in\mathcal{L}_{B\ominus L_{1}\ominus\dots\ominus L_{k^{\prime}}}\}$
where $V_{0}(L_{1},\dots,L_{k}):=\text{Id}$. The natural filter algorithm˜1 for $k$ queries in $\mathcal{L}$ with budget $B$ is free if and only if $V_{k}(\lambda)\preceq B$.
Informally, $V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{k^{\prime}}$. Naturally, when the adversary has already selected all of its moves, the remaining privacy cost is Id, i.e. nothing.
It is also of interest to understand when a given family of queries $\mathcal{L}$ simultaneously admits free natural filters against an arbitrary budget $B$. We show that this is the case exactly when the supremum commutes with the convolution operator.
@@
Fix a family of PLDs $\mathcal{L}$. The natural filter Algorithm˜1 is free for every budget $B$ simultaneously and query capacity $k$ if and only if $\mathcal{L}$ has the following property:
$\forall L\in\mathcal{L}\ \textrm{and}\ \mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\quad L\oplus\sup{\mathcal{L}^{\prime}}=\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}.$
@@
First, assume that the natural filter is free for every budget $B$ and for any query capacity. Now, let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$. For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$. As for the reverse inequality, consider the rule
$\Gamma(\lambda):=\{L\}\quad\quad\Gamma(L):=\mathcal{L}^{\prime}$
@@
of length $2$ and the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$. Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$. Indeed, let $M=M_{1}\oplus M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that we have that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k-1}\in\mathcal{L}^{<k}$. For each $L_{i}$, we can find a mechanism $M^{i}$ with $\operatorname{PLD}(M^{i})=L_{i}$. Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M^{1}\otimes\dots\otimes M^{k-1}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k^{\prime}}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$. Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup\{L\oplus\mathcal{L}^{\prime}\}.$
On the other hand, assume that convolution commutes with supremum in $\mathcal{L}^{\infty}$. Let $B$ be any budget and let $k$ be a positive integer. With an eye toward Theorem˜1, we claim that the maximal remaining privacy cost of an adversary that has selected some if its moves already satisfies
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})=\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
for $0\leq k^{\prime}<k$. Indeed,
$V_{1}(L_{1},\dots,L_{k-1})$ $=\sup\{L_{k}\oplus V_{0}(L_{1},\dots,L_{k}):L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k}:L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and, if the claim holds for $k^{\prime}$, then
$V$ $_{k-(k^{\prime}-1)}(L_{1},\dots,L_{k^{\prime}-1})$
$=\sup\{L_{k^{\prime}}\oplus V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{\sup\{L_{k^{\prime}}\oplus L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\dots\oplus L_{k}:L_{k^{\prime}},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and the claim holds by induction. In particular,
$V_{k}(\lambda)=\sup\{L_{1}\oplus\dots\oplus L_{k}:L_{1},\dots,L_{k}\in\mathcal{L},L_{1}\oplus\dots\oplus L_{k}\preceq B\}\preceq B$
and so the natural filter is free for $k$ queries in $\mathcal{L}$ with budget $B$ by Theorem˜1. ∎
Our key result is that that this commutativity condition can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{\infty}$ is well-ordered.

### v2

@@ 3 Query-Restricted Adversaries
Taking suprema, we get $\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}\otimes M^{\star}}$ and thus $\operatorname{PLD}(\Gamma)\succeq L\oplus\operatorname{PLD}(\Gamma|_{L})$ by Propositions˜11 and 3, which completes the proof since $L$ was arbitrary. ∎
A final tool we will require for analyzing adversaries respecting a privacy rule is that privacy rules of infinite length can be understood by passing to limits of finite privacy rules.
Lemma 2.
Let $\Gamma$ be a privacy rule of infinite length. For $k\in\mathbb{Z}^{+}$, let $\Gamma^{k}$ be the privacy rule of length $k$ denoting the restriction of $\Gamma$ to query sequences of length less than $k$. Then
$\operatorname{PLD}(\Gamma)=\sup\{\operatorname{PLD}(\Gamma^{k}):k\in\mathbb{Z}^{+}\}.$
Proof.
It is immediate that $\operatorname{PLD}(\Gamma^{k})\preceq\operatorname{PLD}(\Gamma)$ for all $k\in\mathbb{Z}^{+}$ because any mechanism $M\in\operatorname{Adv}(\Gamma^{k})$ can be extended to a mechanism in $\operatorname{Adv}(\Gamma)$ with the same privacy loss by simply repeatedly issuing the Id query after the first $k$ rounds.
As for the reverse inequality, we pass again to the hockey-stick curve. To this end, let $M\in\operatorname{Adv}(\Gamma)$, let $\ell$ denote the likelihood ratio for $M$, let $M^{k}$ denote the restriction of $M$ to the first $k$ outputs, and let $\ell^{k}$ denote the likelihood ratio of $M^{k}$. Denoting $y^{k}:=(y_{1},\dots,y_{k})$ for $y=(y_{1},y_{2},\dots)$, it is clear that $\ell(y)=\lim_{k\to\infty}\ell^{k}(y^{k})$. By Proposition˜5 and the dominated convergence theorem, we have
$h_{M}(x)$ $=\mathop{\mathbb{E}}_{Y\sim M(D)}[(1-x/\ell(Y))_{+}]$
$=\mathop{\mathbb{E}}_{Y\sim M(D)}[\lim_{k\to\infty}(1-x/\ell^{k}(Y^{k}))_{+}]$
$=\lim_{k\to\infty}\mathop{\mathbb{E}}_{Y\sim M(D)}[(1-x/\ell^{k}(Y^{k}))_{+}]$
$=\lim_{k\to\infty}\mathop{\mathbb{E}}_{Y\sim M^{k}(D)}[(1-x/\ell^{k}(Y))_{+}]$
$=\lim_{k\to\infty}h_{M^{k}}(x)$
$\leq\sup\{h_{M^{k}}(x):k\in\mathbb{Z}^{+}\}.$
Since clearly $M^{k}\in\operatorname{Adv}(\Gamma^{k})$, the desired inequality follows from Proposition˜11. ∎
4 Residue Filters
In this section, we develop a new class of privacy filters capable of leveraging exact privacy accounting: the residue filters. We show that residue filters capture all of the filters available in the literature including the pure DP filter, the Rényi filter, the GDP filter, as well as the natural filter.
Algorithm 1 UpdateFilter
0: Analyst $\mathcal{A}$, update rule $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$, budget $B\in\mathcal{B}$, dataset $D$
for $i\leftarrow 1,2,\dots$ do
  $\mathcal{A}(Y_{1},\dots,Y_{i-1})$ gives mechanism $M_{i}$ that has PLD $L\in\mathcal{L}$
  if $L\preceq B$ then
   yield $Y_{i}\sim M_{i}(D)$
   $B\leftarrow U(B,L)$
  else
   break
A key piece of structure that appears to enable existing privacy filters is the notion of a remaining privacy budget. This same filter structure can be outfitted naturally with exact accounting. For this, we require some simple ingredients: Let $\mathcal{B}$ be a set of PLDs representing legal budgets, let $\mathcal{L}$ be a family of allowed queries, and let $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$ be a map taking a legal budget and query and producing an updated budget. For simplicity, we assume $\text{Id}\in\mathcal{B},\mathcal{L}$. Algorithm˜1 formalizes the interaction between analyst and the update filter with initial budget $B$. Now, notice that Algorithm˜1 can be captured succinctly by the recursive privacy rule
$\operatorname{Update}_{U,B}(\lambda)$ $:=\{L\in\mathcal{L}:L\preceq B\}$
$\operatorname{Update}_{U,B}|_{L_{1}}$ $:=\textrm{Update}_{U,U(B,L_{1})}.$
Our key result is that privacy filters derived from budget updates are valid as long as the updates satisfy a natural residue condition. Informally, it is enough that every budget encountered during the interaction dominates the privacy cost of the next budget plus the cost of the next query. In this case, we say that the updated budget is a residue of the prior budget with respect to the query.
Theorem 1.
Fix an update rule $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$ and suppose that, for every budget $B\in\mathcal{B}$ and every $L\in\mathcal{L}$ with $L\preceq B$, we have
$U(B,L)\oplus L\preceq B.$ ($\star$)
Then, for any initial budget $B\in\mathcal{B}$, the update filter Algorithm˜1 is free, i.e. $\operatorname{PLD}(\operatorname{Update}_{U,B})\preceq B$. In this case, we call the filter a residue filter.
Proof.
By Lemma˜2, it suffices to prove that the finite rule $\operatorname{Update}_{U,B}^{k}$ is free for all $B\in\mathcal{B}$ and $k\in\mathbb{Z}^{+}$ under the residue condition. We proceed by induction. For $k=1$, we have
$\operatorname{PLD}(\operatorname{Update}_{U,B}^{k})=\sup\{L\in\mathcal{L}:L\preceq B\}\preceq B$
for any $B\in\mathcal{B}$. Moreover, if this is true for $k\in\mathbb{Z}^{+}$, then by Lemma˜1 and by assumption,
$\operatorname{PLD}(\operatorname{Update}_{U,B}^{k+1})$ $=\sup\{L\oplus\operatorname{PLD}(\operatorname{Update}_{U,B}^{k+1}|_{L}):L\in\mathcal{L},L\preceq B\}$
$=\sup\{L\oplus\operatorname{PLD}(\operatorname{Update}_{U,U(B,L)}^{k}):L\in\mathcal{L},L\preceq B\}$
$\preceq\sup\{L\oplus U(B,L):L\in\mathcal{L},L\preceq B\}\preceq B$
for any $B\in\mathcal{B}$ as well. ∎
It turns out that residue filters already capture the existing privacy filters in addition to the natural filter, as we will see.
Remark 2.
The GDP filter Koskela et al. (2022), which permits a sequence of queries that are $\mu_{1},\dots,\mu_{t}$-GDP as long as $\mu_{1}^{2}+\dots+\mu_{t}^{2}\leq\mu^{2}$ for a fixed budget $\mu$, is exactly the residue filter with initial budget $B=G_{\mu}$ and update rule $U(G_{\mu},L):=G_{\sqrt{\mu^{2}-\nu^{2}}}$ where $\nu\geq 0$ is minimal such that $L\preceq G_{\nu}$. Because
$U(G_{\mu},L)+L\preceq G_{\sqrt{\mu^{2}-\nu^{2}}}+G_{\nu}=G_{\mu},$
Theorem˜1 immediately yields that the GDP filter is free.
Remark 3.
The pure DP filter Rogers et al. (2016), which permits a sequence of queries that are $\varepsilon_{1},\dots,\varepsilon_{t}$-DP as long as $\varepsilon_{1}+\dots+\varepsilon_{t}\leq\varepsilon$ for a fixed budget $\varepsilon$, is exactly the residue filter with initial budget $B=R_{\varepsilon}$ and update rule $U(R_{\varepsilon},L):=R_{\varepsilon-\varepsilon_{L}}$ where $\varepsilon_{L}\geq 0$ is minimal such that $L\preceq R_{\varepsilon_{L}}$. Because
$U(R_{\varepsilon},L)+L\preceq R_{\varepsilon-\varepsilon_{L}}+R_{\varepsilon_{L}}\preceq R_{\varepsilon},$
Theorem˜1 immediately yields that the pure DP filter is free.
Unlike the GDP and pure DP forms of differential privacy, Rényi DP Feldman and Zrnic (2021) cannot be certified exactly by domination by a canonical PLD, so the Rényi filter cannot directly be expressed as a residue filter. Nonetheless:
Remark 4.
The Rényi filter Feldman and Zrnic (2021), which permits a sequence of queries that are $(\alpha,\rho_{1}),\dots,(\alpha,\rho_{t})$-RDP as long as $\rho_{1}+\dots+\rho_{t}\leq\rho$ for a fixed budget $\rho$ and order $\alpha$, can be expressed as a budget maintaining filter satisfying a residue condition. The maintained budget is initially $\rho$ and, upon receiving an $(\alpha,\rho_{i})$-RDP query, the filter accepts the query as long as $\rho_{i}\leq\rho$ and updates the budget to $\rho-\rho_{i}$. By linearity properties of RDP, this update satisfies an analogous residue condition to ($\star$1) and hence the same techniques can be used to show that this filter is free.
In particular, derivative filters including the zCDP and the advanced composition filter seen in e.g. Whitehouse et al. (2023) are also a form of residue filter.
Perhaps more notable is that the natural filter, whenever it is free, can also be understood as a residue filter. We will discuss this in detail in §5. The key takeaway is that residue filters either directly or in essence subsume previous privacy filters as well as the natural filter itself. This is strongly suggestive evidence that filters based on update rules satisfying a residue-like condition should be viewed as a standard technique for constructing new privacy filters. We strengthen this claim by constructing a new filter for GDP budgets that strictly improves upon the GDP filter.
Theorem 5.
Consider the update rule
$U(G_{\mu},L):=G_{\mu^{\prime}},\quad\mu^{\prime}\geq 0\text{ maximal such that }G_{\mu^{\prime}}\oplus L\preceq G_{\mu}$
for GDP budgets $G_{\mu}$ and arbitrary queries $L\preceq G_{\mu}$. This rule satisfies the residue condition ($\star$1) and hence the update filter derived from it is a residue filter. We call this filter the GDP residue filter.
(a) Nearly Gaussian query: gain of $\mu\approx 0.036$ GDP budget.
(b) Nearly Gaussian query, low budget: gain of $\mu\approx 0.078$ GDP budget.
(c) Non-Gaussian query: gain of $\mu\approx 0.147$ GDP budget.
Figure 1: A naïve GDP filter directly approximates the query by GDP and determines a new budget $\mu_{naive}$. When composing $G_{\mu_{naive}}$ with the query (purple), a gap with the original budget shows that budget was wasted. The GDP residue filter instead searches for the largest valid residual GDP budget $\mu_{res}$. Composing $G_{\mu_{res}}$ with the query almost matches the budget: very little is wasted.
It is immediate that this update rule satisfies the residue condition [^1] One can show carefully that the maximum exists by noticing that $\mathcal{L}_{\mathrm{gdp}}$ is totally-ordered and applying Lemma 3, which we develop in the next section.. The update can be computed efficiently by binary search over $\mu^{\prime}\in[0,\mu]$. It is also clear that the GDP residue filter cannot produce an updated budget smaller than the standard GDP filter. What is perhaps less obvious is that the GDP residue filter is different from the GDP filter. Indeed, intuitively the GDP filter may significantly overestimate the cost of a query $L\preceq G_{\nu}$ if $L$ is very hard to approximate by $G_{\nu}$. On the other hand, even if $L$ itself is badly approximated by $G_{\nu}$, the residue class $\{B^{\prime}:B^{\prime}\oplus L\preceq G_{\mu}\}$ that remains when removing $L$ from $G_{\mu}$ may admit a much tighter approximation.
Figure˜1 illustrates using $f$-DP curves the savings of our residual filter in three different privacy regimes. Each plot shows an initial GDP budget, a pure DP query, the updated budget under the naïve GDP filter update ($\mu_{naive}$), and the updated budget under our GDP residue filter $\mu_{res}$. Generally, the residue filter always outperforms the naïve filter: $\mu_{res}>\mu_{naive}$. When the query is well-approximated by GDP and the budget is much larger than the cost of the query, the gains are fairly small, though still meaningful (fig.˜1(a)). When the query is well-approximated by GDP but the budget is small, perhaps due to depletion from previous queries, the gains are much more significant (fig.˜1(b)). The gains are predictably most significant for queries that are poorly approximated by GDP (fig.˜1(c)). In each case, the residual GDP budget, when composed with the query, is very close to the starting budget, showing that most of the budget has in fact been put to use.
@@ 5 Characterizing the Natural Filter
We now turn toward the natural filter Algorithm˜2 with some budget $B$ and queries restricted to some set $\mathcal{L}$. Our goal is to understand when this filter comes at no additional privacy cost, i.e. the mechanism induced by the interaction between the filter and an adversarial analyst itself has privacy bounded by $B$. In this case, we say that the natural filter is free. For simplicity, we will assume that $\text{Id}\in\mathcal{L}$.
@@
0: Budget $B$, analyst $\mathcal{A}$, legal query set $\mathcal{L}$, query capacity $k\in\mathbb{Z}^{+}\cup\{\infty\}$, dataset $D$
for $t\leftarrow 1,\dots,k$ do
  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}$
  if $L_{1}\oplus\dots\oplus L_{t}\preceq B$ then
   yield $Y_{t}\sim M_{t}(D)$
  else
   break
Now, notice that the natural filter Algorithm˜2 with budget $B$, query set $\mathcal{L}$, and capacity $k$ is free exactly when $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$ where
$\operatorname{Nat}_{\mathcal{L},B}(L_{1},\dots,L_{t}):=\{L_{t+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{t+1}\preceq B\}.$
In particular, by applying Lemma˜1 to the rule $\operatorname{Nat}_{\mathcal{L},B}^{k}$, we immediately get the following characterization of the cost of interacting with the natural filter.
@@
Let $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$, let $B\in\mathcal{L}_{\mathrm{all}}$, and consider, for $L^{\prime}\preceq B$, the PLD operator
$V_{k+1}(L^{\prime}):=\sup\{L\oplus V_{k}(L^{\prime}\oplus L):L\in\mathcal{L},L^{\prime}\oplus L\preceq B\}$
where $V_{1}(L^{\prime}):=\sup\{L\in\mathcal{L}:L^{\prime}\oplus L\preceq B\}$. Then, for any $k\in\mathbb{Z}^{+}$ and legal $L_{1},\dots,L_{t}\in\mathcal{L}$, we have $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k}\mid_{L_{1},\dots,L_{t}})=V_{k-t}(L_{1}\oplus\dots\oplus L_{t})$ and, in particular, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})=V_{k}(\text{Id})$.
Informally, $V_{k-t}(L_{1}\oplus\dots\oplus L_{t})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{t}$. Naturally, when only one move is left to select, the effect of adaptivity vanishes and the adversary simply selects the valid move of maximal cost.
It is also of interest to understand when a given family of queries $\mathcal{L}$ admits free natural filters against an arbitrary budget $B$. When a filter is free against all budgets, we say that the filter is universally free. In the case of the natural filter, we show that it is universally free exactly when it can be realized as a residue filter. For convenience, given $k\in\mathbb{Z}^{+}\cup\{\infty\}$ we let $\mathcal{L}^{<k}:=\{L_{1}\oplus\dots\oplus L_{t}:L_{1},\dots,L_{t}\in\mathcal{L},0\leq t<k\}$ denote the closure of $\mathcal{L}$ under fewer than $k$ convolutions and we let $L\oplus\mathcal{L}^{\prime}:=\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$ for a set of PLDs $\mathcal{L}^{\prime}$.
@@
Fix $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$, $k\in\mathbb{Z}^{+}\cup\{\infty\}$, and consider the update rule $U:\mathcal{L}_{\mathrm{all}}\times\mathcal{L}\to\mathcal{L}_{\mathrm{all}}$ given by $U(B,L):=\sup\{L^{\prime}\in\mathcal{L}^{<k}:L\oplus L^{\prime}\preceq B\}$. The following are equivalent:
  1. (i)
    For all $B\in\mathcal{L}_{\mathrm{all}}$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$;
  2. (ii)
    For all $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$, $L\oplus\sup{\mathcal{L}^{\prime}}=\sup(L\oplus\mathcal{L}^{\prime})$;
  3. (iii)
    For all $B\in\mathcal{L}_{\mathrm{all}}$ and $L\in\mathcal{L}$, $U(B,L)\oplus L\preceq B$; and
  4. (iv)
    For all $B\in\mathcal{L}_{\mathrm{all}}$, $\operatorname{PLD}(\operatorname{Update}_{U,B}^{k})\preceq B$.
In any of these cases, the natural filter $\operatorname{Nat}_{\mathcal{L},B}^{k}$ and the residue filter $\operatorname{Update}_{U,B}^{k}$ are identical.
@@
Without loss of generality, by Lemma˜2, assume that $k$ is finite.
@@
First, we show that (i) implies (ii), so suppose that the natural filter is free for every budget $B$ and let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$. For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup(L\oplus\mathcal{L}^{\prime})\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free. As for the reverse inequality, consider the length-$2$ rule $\Gamma(\lambda):=\{L\},\Gamma(L):=\mathcal{L}^{\prime}$ as well as the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$. Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$. Indeed, let $M=M_{1}\otimes M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{2,1}\oplus\dots\oplus L_{2,t}\in\mathcal{L}^{<k}$. For each $L_{2,i}$, we can find a mechanism $M_{2,i}$ with $\operatorname{PLD}(M_{2,i})=L_{2,i}$. Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M_{2,1}\otimes\dots\otimes M_{2,t}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{2,1}\oplus\dots\oplus L_{2,t}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$. Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup(L\oplus\mathcal{L}^{\prime}).$
That (ii) implies (iii) is immediate from the construction of the update rule $U$ by taking $\mathcal{L}^{\prime}:=\{L^{\prime}\in\mathcal{L}^{<k}:L\oplus L^{\prime}\preceq B\}$; likewise (iii) implies (iv) follows immediately from Theorem˜1, so what remains to show is that, if the residue filter $\operatorname{Update}_{U,B}$ is universally free, then so is $\operatorname{Nat}_{\mathcal{L},B}$, in which case the two filters coincide.
To this end, first notice that $\operatorname{Update}_{U,B}^{k}$ is always at least as permissive a filter as $\operatorname{Nat}_{\mathcal{L},B}^{k}$. Indeed, for any $L\in\mathcal{L}$ and $L^{\prime}\in\mathcal{L}^{<\infty}$ such that $L\oplus L^{\prime}\preceq B$, it follows immediately from the construction of $U$ that $L^{\prime}\preceq U(B,L)$. Thus for any legal query sequence $L_{1},\dots,L_{t}$,
$L_{t+1}\in\operatorname{Nat}_{\mathcal{L},B}^{k}(L_{1},\dots,L_{t})$ $\implies L_{1}\oplus\dots\oplus L_{t+1}\preceq B$
$\implies L_{2}\oplus\dots\oplus L_{t+1}\preceq U(B,L_{1})$
$\implies\dots$
$\implies L_{t+1}\preceq U(\dots U(B,L_{1})\dots,L_{t})$
$\implies L_{t+1}\in\operatorname{Update}_{U,B}^{k}(L_{1},\dots,L_{t}).$
Therefore, if $\operatorname{Update}_{U,B}^{k}$ is universally free, then so is $\operatorname{Nat}_{\mathcal{L},B}^{k}$.
Finally, notice that, as long as the residue condition for $U$ is satisfied, then $\operatorname{Nat}_{\mathcal{L},B}^{k}$ is at least as permissive a filter as $\operatorname{Update}_{U,B}^{k}$. Indeed, for any $L\in\mathcal{L}$ and $L^{\prime}\in\mathcal{L}^{<\infty}$ for which $L^{\prime}\preceq U(B,L)$, Proposition˜2 and the residue condition ensure that $L\oplus L^{\prime}\preceq L\oplus U(B,L)\preceq B$. Inductively, just as before, this implies that for any legal query sequence $L_{1},\dots,L_{t}$, we have that $L_{t+1}\in\operatorname{Update}(U,B)^{k}(L_{1},\dots,L_{t})\implies L_{t+1}\in\operatorname{Nat}_{\mathcal{L},B}^{k}(L_{1},\dots,L_{t})$ and so the two filters are indeed equivalent. ∎
Our key result is that that these conditions can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{<k}$ is totally-ordered. For a distribution $A$ over $\overline{\mathbb{R}}$, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the finite points with locally positive probability mass. We say that $A$ is non-degenerate if $|\operatorname{supp}(A)|>1$, i.e. $A$ is supported at at least two finite points.